OpenStudy (mayankdevnani):

Two cars A and B are moving towards each other on a horizontal surface. The car A has mass 60 g and moves towards the right with speed 60 cm/s. The car B has a mass of 100 g and moves towards the left with a speed of 20 m/s. The two cars collide and get stuck to each other. With what velocity will they move after the collision?

OpenStudy (mayankdevnani):

@mathslover @igbasallote @ghazi

OpenStudy (mayankdevnani):

@sauravshakya @UnkleRhaukus @ganeshie8

OpenStudy (mayankdevnani):

@Ishaan94

OpenStudy (ghazi):

use conservation of momentum

OpenStudy (mayankdevnani):

@hartnn

OpenStudy (mayankdevnani):

ok!! then

OpenStudy (ghazi):

write the equation for conservation of momentum first

OpenStudy (mayankdevnani):

momentum before collision = momentum after collision

OpenStudy (ghazi):

m1*v1+m2v2= m1vf+m2vf

OpenStudy (ghazi):

that is m1v1+m2v2=(m1+m2)vf

OpenStudy (mayankdevnani):

no, but m1*u1+m2+u2=m1v1+m2v2 @ghazi

OpenStudy (ghazi):

\[0.060*0.6+0.1100*20=(60+100)*V _{f}\] dude go for the equation of perfectly inelastic collision ...and then solve this problem

OpenStudy (ghazi):

didn't you see in your question...that two body were moving together

OpenStudy (mayankdevnani):

sry...

OpenStudy (mayankdevnani):

then

OpenStudy (ghazi):

just solve the above equation

OpenStudy (mayankdevnani):

@ghazi answer is in cm/s

OpenStudy (mayankdevnani):

plz solve this eq. plz @ghazi plz.....

OpenStudy (ghazi):

okayy..convert that into cm

OpenStudy (mayankdevnani):

yaa..

OpenStudy (mayankdevnani):

plz solve

OpenStudy (mayankdevnani):

ok thanx...

OpenStudy (mayankdevnani):

@ghazi thnx fr help

OpenStudy (ghazi):

\[(0.060*60+0.100*200)= (100)*V _{f}.....V _{f}=\frac{ 3.60+20 }{ 0.100 }=236 cm/s\]

OpenStudy (mayankdevnani):

no, answer is 10 cm/s

OpenStudy (ghazi):

there must be some miscalculation but this is the equation that will be used...i might have done some wrong conversion ...

OpenStudy (mayankdevnani):

ok

mathslover (mathslover):

m_A v_A + m_b B v_B = Momentum after collision